# Feedback-Enhanced Online Multiple Testing with Applications to Conformal Selection

### v1

@@ Abstract
We study online multiple testing with feedback, where decisions are made sequentially and the true state of the hypothesis is revealed after the decision has been made, either instantly or with a delay. We propose GAIF, a feedback-enhanced generalized alpha-investing framework that dynamically adjusts thresholds using revealed outcomes, ensuring finite-sample false discovery rate (FDR)/marginal FDR control. Extending GAIF to online conformal testing, we construct independent conformal $p$-values and introduce a feedback-driven model selection criterion to identify the best model/score, thereby improving statistical power. We demonstrate the effectiveness of our methods through numerical simulations and real-data applications.
@@
Keywords: Conformal prediction; Distributional shifts; Generalized alpha-investing procedure; Model selection; Online conformal $p$-value; Online FDR control
@@ 1 Introduction
Real-time decision making plays a critical role in a growing number of modern applications, such as online recruitment for job hiring [Faliagka et al., 2014], real-time alignment of large language models [Huang et al., 2025], and time-series anomaly detection [Rebjock et al., 2021], etc. These tasks can be naturally formulated as online multiple testing problems [Foster and Stine, 2008]. Consider a potentially infinite stream of null hypotheses $\{\mathbb{H}_{01},\mathbb{H}_{02},\dots,\mathbb{H}_{0t},\dots\}$ tested sequentially based on arriving test statistics $\{z_{1},z_{2},\dots\}$. At each time $t$, a real-time decision must be made upon observing $z_{t}$. Let $\theta_{t}$ denote the true state of the hypothesis at time $t$, where $\theta_{t}=0$ if $\mathbb{H}_{0t}$ is true and $\theta_{t}=1$ otherwise. The testing problem at time $t$ is
@@
Unlike the classical setting, we consider another realistic scenario in which the decision-maker receives real-time feedback on $\theta_{t}$ after issuing the decision $\delta_{t}$. Concretely, at each time $t$, the decision $\delta_{t}$ is based on the current statistic $z_{t}$, the history $\{\delta_{1},\dots,\delta_{t-1}\}$, and any feedback observed up to that point. The feedback $\theta_{t}$ may arrive immediately or with some delay (for example, as $\theta_{t-d}$ for $d\geq 1$) before moving on to time $t+1$. This feedback-available framework naturally arises in many practical applications. To demonstrate its relevance, we present three motivating examples: online conformal selection, real-time alignment of large language models, and time-series anomaly detection.
  • Online conformal selection. Conformal selection aims to identify valuable individuals whose unknown label $Y$ satisfies a pre-specified requirement by leveraging machine learning predictions [Jin and Candès, 2023a, Wu et al., 2024]. The online conformal selection setting naturally aligns with the feedback-available online multiple testing framework. For example, in diabetes risk prediction, a patient’s true condition may later be confirmed by an expert, offering feedback that could improve future decisions. Some works [Huo et al., 2024, Xu and Ramdas, 2024] considered online conformal testing problems but did not exploit feedback information.
@@
  • Time series anomaly detection. Detecting anomalies in time series is crucial for industrial monitoring, fraud detection, and healthcare analytics. To ensure reliability, prior work [Gang et al., 2021, Rebjock et al., 2021, Krönert et al., 2023] addresses online FDR control but typically ignores real-time feedback. In practice, such feedback is often available: once an anomaly is flagged, a subsequent system failure, user verification of a fraudulent transaction, or expert annotation may confirm or refute the alarm.
These observations underscore that real-time feedback is a foundational element of adaptive decision-making, playing a pivotal role across a wide range of online applications. Despite its importance, current online multiple testing procedures seldom incorporate feedback in a systematic way. This naturally raises a fundamental and compelling question: Can real-time feedback be effectively incorporated into online multiple testing procedures in a way that allows us to enhance statistical power while still ensuring valid error rate control? To this end, we develop a new framework that systematically integrates feedback into the online testing process, achieving significant performance gains without compromising statistical validity. Specifically, our proposed strategies utilize feedback through three key mechanisms: (1) within the proposed GAIF framework—a feedback-enhanced extension of generalized alpha-investing (GAI) [Ramdas et al., 2017, Javanmard and Montanari, 2018]—we adaptively adjust testing thresholds based on past decisions and revealed outcomes together; (2) by updating an online calibration dataset—used to generate explicit, valid, and independent online $p$-values for conformal testing; and (3) by guiding model selection in online conformal testing. These enhancements collectively improve power over existing GAI methods to a large degree, while still ensuring valid online FDR control.
@@ 1.1 Our contributions
To the best of our knowledge, this is the first work to incorporate feedback information directly into the construction of testing thresholds for online FDR procedures and to extend this idea to the setting of online conformal testing. Our main contributions are twofold:
  • Generalized alpha-investing with feedback (GAIF). GAIF is both comprehensive and flexible: it can enhance almost any existing method within the traditional GAI family, and it naturally extends to more complex scenarios—such as delayed feedback and local dependence. We prove that GAIF-based procedures maintain valid FDR control under independence and valid mFDR control under local dependence.
  • Online conformal multiple testing. We extend GAIF to the setting of online conformal multiple testing, providing explicit construction of valid, independent $p$-values (for null hypotheses) by dynamically updating the calibration dataset in this context. By suitably modifying the GAIF rules, we obtain a finite-sample guarantee of mFDR control for our procedure, online conformal testing with feedback (OCTF). This extension bridges the gap between traditional online multiple testing and conformal inference, yielding distribution-free, model-agnostic tools for real-time decision-making. Moreover, we introduce an online model-selection criterion: predictive models and conformity scores are chosen adaptively based on feedback, further improving power, especially under distribution shifts among non-nulls.
We provide rigorous proofs for all proposed methods. Extensive simulations and real-data experiments demonstrate that our procedures substantially outperform existing approaches while effectively controlling the online FDR when feedback information is provided.
@@ 1.2 Related works
Early works on online multiple testing began with the alpha-investing strategy of Foster and Stine [2008], later generalized by Aharoni and Rosset [2014] and Javanmard and Montanari [2018] into the generalized alpha-investing (GAI) framework, which led to the LORD algorithm. Building on this line of work, Ramdas et al. [2017] introduced LORD++, an improved version of GAI tailored for online FDR control. Subsequent refinements include SAFFRON [Ramdas et al., 2018], which adapts to the proportion of non-nulls, and ADDIS [Tian and Ramdas, 2019], which adjusts for conservative null $p$-values. These methods, including LORD++, SAFFRON, and ADDIS, achieve online FDR control when null $p$-values are independent of all other $p$-values. Separately, Gang et al. [2021] developed structure-adaptive rules based on local FDR, which improve power but only ensure asymptotic FDR control under correct model specification. For a comprehensive review, see Robertson et al. [2023]. All the above methods determine thresholds solely from past rejections, without considering real-time feedback.
@@
In practice, hypotheses often exhibit dependence, and applying methods designed for independence can lead to inflated error rates. To address arbitrary dependence, Xu and Ramdas [2024] proposed e-LOND, an FDR-controlling procedure based on e-values. Zhang et al. [2025] extended this approach to e-GAI, achieving improved power by dynamically allocating the testing levels. Alternatively, research has focused on local dependence structures. Zrnic et al. [2021] introduced LORDdep and SAFFRONdep, establishing mFDR control under local dependence; Rebjock et al. [2021] later adapted these methods to time-series anomaly detection. Recently, Fisher [2024] showed that LORD++ with suitable local modifications can maintain FDR control under certain dependence, while Fischer et al. [2024] proposed an online procedure under PRDS dependence. Despite these advances, existing dependence-aware methods do not incorporate any real-time feedback.
@@
Conformal inference [Vovk et al., 2005] offers a model-agnostic way to quantify prediction uncertainty. In the multiple testing setting, early works constructed conformal $p$-values and applied the Benjamini–Hochberg (BH) procedure [Benjamini and Hochberg, 1995] to achieve finite-sample FDR control [Bates et al., 2023, Jin and Candès, 2023a]. Subsequent extensions addressed covariate shift [Jin and Candès, 2023b], constrained selection [Wu et al., 2024, Nair et al., 2025], and conditional testing [Wu et al., 2025], as well as model selection [Bai and Jin, 2024, Gui et al., 2025]. However, these contributions remain confined to offline settings. Few efforts to extend conformal multiple testing to the online domain [Huo et al., 2024, Xu and Ramdas, 2024] have so far overlooked the role of feedback information. Although related research on the construction of conformal prediction sets [Gibbs and Candès, 2021, 2024] has considered feedback, multiple testing problems with FDR control remain largely unexplored. One very recent work, by Humbert et al. [2025], establishes asymptotic online FDP control through an online learning strategy, while we achieve feedback-enhanced testing based on the GAI framework.
@@ 1.3 Organization of the paper
The remainder of this paper is organized as follows. Section 2 introduces the GAIF procedure and establishes FDR control under independence and mFDR control under local dependence. In Section 3, we construct explicit online conformal $p$-values and apply modified GAIF rules to online conformal testing, providing finite-sample theoretical guarantees. We also extend the proposed framework to achieve model selection and address distribution shifts. Simulation and real-data experiment results are presented in Sections 4 and 5, respectively. Finally, we conclude the paper in Section 6.
@@ 2 Generalized Alpha-Investing with Feedback (GAIF)
In this section, we first revisit the traditional generalized alpha-investing (GAI) framework and its extensions in Section 2.1. Then, in Section 2.2, we introduce a new framework named GAIF, that incorporates feedback information. We provide two concrete approaches of GAIF in Section 2.3. In Section 2.4, we establish the finite-sample online FDR control of GAIF procedures under independence. Additionally, we discuss the extensions to address local dependence in Section 2.5.
@@ 2.1 Recap: Generalized alpha-investing
The first approaches for controlling error rates in an online setting were based on so-called “alpha-investing” strategies [Foster and Stine, 2008], and subsequently generalized into the generalized alpha-investing (GAI) framework [Aharoni and Rosset, 2014, Javanmard and Montanari, 2018]. The key idea is to compare each incoming $p$-value $p_{t}$ against a dynamically chosen threshold $\alpha_{t}$, to accumulate additional $\alpha$-wealth” upon each rejection, and to make testing decisions according to $\delta_{t}=\mathbb{I}\{p_{t}\leq\alpha_{t}\}$. Based on GAI, Ramdas et al. [2017] proposed controlling the online FDR by ensuring that an estimate of the FDP remains below a pre-specified level $\alpha$. A specific example is LORD++ algorithm, where the estimated $\operatorname{FDP}$ at time $t$ is given by
$\widehat{\operatorname{FDP}}_{\text{LORD}}(t)=\frac{\sum_{j\leq t}\alpha_{j}}{\sum_{j\leq t}\delta_{j}\vee 1}\overset{(i)}{\geq}\operatorname{FDP}^{*}(t)=\frac{\sum_{j\leq t,j\in\mathcal{H}_{0}}\alpha_{j}}{\sum_{j\leq t}\delta_{j}\vee 1}\approx\operatorname{FDP}(t):=\frac{\sum_{j\leq t,j\in\mathcal{H}_{0}}\mathbb{I}\{p_{j}\leq\alpha_{j}\}}{\sum_{j\leq t}\delta_{j}\vee 1},$ (1)
where $\mathcal{H}_{0}$ is the index set of null hypotheses. To ensure that $\operatorname{FDR}(t)\leq\alpha$, it suffices to enforce $\widehat{\operatorname{FDP}}_{\text{LORD}}(t)=\frac{\sum_{j\leq t}\alpha_{j}}{R(t)\vee 1}\leq\alpha.$ Building on this, some adaptive versions of LORD++ were proposed subsequently [Ramdas et al., 2018, Tian and Ramdas, 2019]. Specifically, similar to Storey-BH [Storey et al., 2004], Ramdas et al. [2018] proposed SAFFRON with a user-specified parameter $\lambda\in[0,1]$ to estimate the null proportion $\pi_{0}$, giving the estimated FDP as
$\widehat{\operatorname{FDP}}_{\text{SAFFRON}}(t)=\frac{\sum_{j\leq t}\alpha_{j}\frac{\mathbb{I}\{p_{j}>\lambda\}}{(1-\lambda)}}{\sum_{j\leq t}\delta_{j}\vee 1}.$
All of the above GAI-based procedures guarantee mFDR control under the conditional super-uniformity assumption for null $p$-values:
@@
where $\mathcal{G}_{t}=\sigma(\delta_{1},\dots,\delta_{t})$ and each threshold $\alpha_{t}=f_{t}(\delta_{1},\dots,\delta_{t-1})$ is measurable with respect to past decisions. Furthermore, if all null $p$-values are independent of other $p$-values, these procedures also ensure strict FDR control.
Note that the inequality (i) in (1) is tight if and only if $\theta_{j}=0$ for all $j\leq t$. This conservative step is one of major sources of gap between the realized FDR of LORD++ and the target level $\alpha$. In our context, since potential feedback becomes available during the testing process, it can serve as a natural remedy.

### v2

@@ Abstract
This work studies online multiple testing with feedback, where decisions are made sequentially, and the true state of the hypothesis is revealed after decisions are made, either instantly or with a delay, and under either full or bandit feedback. We propose Generalized alpha-investing with feedback (GAIF) along with its adaptive variants, a feedback-enhanced framework that dynamically adjusts thresholds using revealed outcomes, ensuring finite-sample false discovery rate (FDR)/marginal FDR (mFDR) control. We further extend GAIF to online conformal testing by constructing valid conformal $p$-values and developing feedback-enhanced testing rules with finite-sample mFDR control. We also propose a feedback-driven score selection criterion to adaptively choose the candidate score that is most effective for the testing procedure, together with a theoretical analysis of its optimality. Numerical simulations and real-data applications demonstrate the effectiveness of our methods.
@@
Keywords: Conformal prediction, distribution shifts, generalized alpha-investing procedure, model selection, online conformal $p$-value, online FDR control
@@ 1 Introduction
Real-time decision making plays a critical role in a growing number of modern applications, such as online recruitment for job hiring [Faliagka et al., 2014], real-time alignment of large language models [Huang et al., 2025], and time-series anomaly detection [Rebjock et al., 2021]. These tasks can be naturally formulated as online multiple testing problems [Foster and Stine, 2008]. Consider a potentially infinite stream of null hypotheses $\{\mathbb{H}_{01},\mathbb{H}_{02},\dots,\mathbb{H}_{0t},\dots\}$ that are tested sequentially based on incoming test statistics such as $p$-values $\{p_{1},p_{2},\dots p_{t},\dots\}$. At each time $t$, a real-time decision must be made based on the current test statistic. Let $\theta_{t}$ denote the true state of the hypothesis at time $t$, where $\theta_{t}=0$ if $\mathbb{H}_{0t}$ is true and $\theta_{t}=1$ otherwise. The testing problem at time $t$ is
@@
Unlike the classical setting, we consider another realistic scenario in which feedback on past hypotheses is revealed after decisions are made. Specifically, after issuing $\delta_{t}$, the corresponding state $\theta_{t}$ may be observed either immediately or after a delay, and in some settings it may be observed only partially. For instance, in the bandit setting [Neu et al., 2010], feedback $\theta_{t}$ is observed only when the decision $\delta_{t}=1$ is made. Our work encompasses all these feedback-available settings, including full or partial feedback, as well as immediate or delayed feedback.
Such feedback-available settings naturally arise in many practical applications. To demonstrate its relevance, we present three motivating examples: online conformal selection, real-time alignment of large language models, and time-series anomaly detection.
  • Online conformal selection. Conformal selection aims to identify valuable individuals whose unknown label $Y$ satisfies a pre-specified requirement by leveraging machine learning predictions [Wu et al., 2024, Jin and Candès, 2025]. The online conformal selection setting naturally aligns with the feedback-available online multiple testing framework. For example, in diabetes risk prediction, a patient’s true condition may later be confirmed by an expert, offering feedback that could improve future decisions. Some works [Huo et al., 2024, Xu and Ramdas, 2024] considered online conformal testing problems but did not exploit feedback information.
@@
  • Time series anomaly detection. Detecting anomalies in time series is crucial for industrial monitoring, fraud detection, and healthcare analytics. To ensure reliability, prior work [Rebjock et al., 2021, Gang et al., 2023, Krönert et al., 2023] addresses online FDR control but typically ignores real-time feedback. In practice, such feedback is often available: once an anomaly is flagged, a subsequent system failure, user verification of a fraudulent transaction, or expert annotation may confirm or refute the alarm.
These observations underscore that feedback is a foundational element of adaptive decision-making, playing a pivotal role across a wide range of online applications. Despite its importance, current online multiple testing procedures seldom incorporate feedback in a systematic way. This naturally raises a fundamental and compelling question: Can available feedback be effectively incorporated into online multiple testing procedures in a way that allows us to enhance statistical power while still ensuring valid error rate control?
To this end, we develop a feedback-enhanced framework that systematically integrates revealed feedback into the generalized alpha-investing (GAI) framework [Aharoni and Rosset, 2014, Javanmard and Montanari, 2018], achieving significant performance gains without compromising statistical validity.
@@ 1.1 Our contributions
To the best of our knowledge, this is the first work to incorporate feedback information directly into the construction of testing thresholds for online FDR procedures and to employ this idea to online conformal testing. Our main contributions are summarized as follows:
  • GAIF and its variants. We introduce Generalized Alpha-Investing with Feedback (GAIF), a feedback-enhanced extension of the GAI framework, together with its adaptive variants. The key idea is to use revealed feedback to reduce the slack in FDP estimation, and, in the adaptive version, to further improve $\alpha$-wealth allocation. The framework applies to a broad range of feedback settings, including full or partial feedback, as well as immediate or delayed feedback, and it also extends to locally dependent scenarios. We establish online FDR control under independence and online mFDR control under local dependence for GAIF and its variants.
  • Online conformal testing with feedback. We extend the GAIF to online conformal testing. We first provide an explicit construction of valid, independent $p$-values (for null hypotheses) by dynamically updating the calibration dataset. We then combine them with suitably modified GAIF rules to obtain online conformal testing with feedback (OCTF), which enjoys finite-sample online mFDR control. This extension bridges the gap between traditional online multiple testing and conformal inference, yielding distribution-free, model-agnostic tools for real-time decision-making.
  • Feedback-driven score selection. We further introduce a strategy to adaptively select among candidate predictive models or conformity scores. This is particularly important under non-stationary non-null distributions, where the optimal score may vary over time. By locally averaging recent non-null $p$-values to track the best-performing score, our selection criterion can accommodate such shifts while maintaining validity.
We provide rigorous proofs for all proposed methods. Extensive simulations and real‐data experiments demonstrate that our procedures substantially outperform existing approaches while effectively controlling the online FDR when feedback information is provided.
@@ 1.2 Related works
Early works on online multiple testing began with the alpha-investing strategy of Foster and Stine [2008], later generalized by Aharoni and Rosset [2014] and Javanmard and Montanari [2018] into the GAI framework, which led to the LORD algorithm. Building on this line of work, Ramdas et al. [2017] introduced LORD++, an improved version of GAI tailored for online FDR control. Subsequent refinements include SAFFRON [Ramdas et al., 2018], which adapts to the proportion of non-nulls, and ADDIS [Tian and Ramdas, 2019], which adjusts for conservative null $p$-values. These methods, including LORD++, SAFFRON, and ADDIS, achieve online FDR control when null $p$-values are independent of all other $p$-values. Separately, Gang et al. [2023] developed structure-adaptive rules based on local FDR, which improve power but only ensure asymptotic FDR control under correct model specification. For a comprehensive review, see Robertson et al. [2023]. All the above methods determine thresholds solely from past rejections, without considering real-time feedback.
@@
In practice, hypotheses often exhibit dependence, and applying methods designed for independence can lead to inflated error rates. To address arbitrary dependence, Xu and Ramdas [2024] proposed e-LOND, an FDR-controlling procedure based on e-values. Zhang et al. [2025] extended this approach to e-GAI, achieving improved power by dynamically allocating the testing levels. Alternatively, research has focused on local dependence structures. Zrnic et al. [2021] introduced LORDdep and SAFFRONdep, establishing mFDR control under local dependence; Rebjock et al. [2021] later adapted these methods to time-series anomaly detection. Recently, Fisher [2024] showed that LORD++ with suitable local modifications can maintain FDR control under certain dependence, while Fischer et al. [2024] proposed an online procedure for p-values satisfying positive regression dependence on a subset (PRDS). Despite these advances, existing dependence-aware methods do not incorporate any real-time feedback.
@@
Conformal inference [Vovk et al., 2005] offers a model-agnostic way to quantify prediction uncertainty. In the multiple testing setting, early works constructed conformal $p$-values and applied the Benjamini–Hochberg (BH) procedure [Benjamini and Hochberg, 1995] to achieve finite-sample FDR control [Bates et al., 2023, Jin and Candès, 2025]. Subsequent extensions addressed covariate shift [Jin and Candès, 2023], constrained selection [Wu et al., 2024, Nair et al., 2025], and conditional testing [Wu et al., 2025], as well as model selection [Bai and Jin, 2024, Gui et al., 2025]. However, these contributions remain confined to offline settings. Few efforts to extend conformal multiple testing to the online domain [Huo et al., 2024, Xu and Ramdas, 2024, Liu et al., 2026] did not take the feedback information into account. Although related research on the construction of online conformal prediction sets [Gibbs and Candès, 2021, 2024, Xi et al., 2025] has considered feedback, multiple testing problems with FDR control remain largely unexplored. One very recent work, by Humbert et al. [2025], establishes asymptotic online FDP control through an online learning strategy, while we achieve feedback-enhanced testing based on the GAI framework.
@@ 1.3 Organization of the paper
The remainder of this paper is organized as follows. Section 2 introduces the GAIF and establishes FDR control under independence. Section 3 extends the GAIF to achieve adaptive $\alpha$-wealth allocation and to address local dependence. In Section 4, we construct explicit online conformal $p$-values and apply modified GAIF rules to online conformal testing, providing finite-sample theoretical guarantees. We also extend the proposed framework to achieve score selection and address distribution shifts. Simulation and real-data experiment results are presented in Sections 5 and 6, respectively. Finally, we conclude the paper in Section 7.
@@ 2 Generalized Alpha-Investing with Feedback (GAIF)
In this section, we begin by revisiting the generalized alpha-investing (GAI) framework in Section 2.1. In Section 2.2, we introduce our feedback-enhanced framework, termed generalized alpha-investing with feedback (GAIF), which incorporates feedback information to improve FDP estimation. In Section 2.3, we establish finite-sample online FDR control for GAIF procedures.
@@ 2.1 Recap: Generalized alpha-investing
Early approaches to error rate control in an online setting were based on so‐called “alpha‐investing” strategies [Foster and Stine, 2008], which were subsequently generalized into the GAI framework [Aharoni and Rosset, 2014, Javanmard and Montanari, 2018]. The key idea is to compare each incoming pp‐value $p_{t}$ against a dynamically chosen threshold $\alpha_{t}$, to accumulate additional $\alpha$‐wealth” upon each rejection, and to make testing decisions according to $\delta_{t}=\mathbb{I}\{p_{t}\leq\alpha_{t}\}$.
Based on GAI, Ramdas et al. [2017] proposed controlling online FDR by ensuring that an estimator of the FDP remains below a pre‐specified level $\alpha$. A specific example is the LORD++ algorithm, where the estimated $\operatorname{FDP}$ at time $t$ is given by
$\widehat{\operatorname{FDP}}_{\text{LORD}}(t)=\frac{\sum_{j\leq t}\alpha_{j}}{1\vee\sum_{j\leq t}\delta_{j}}\overset{(i)}{\geq}\operatorname{FDP}^{*}(t)=\frac{\sum_{j\leq t,j\in\mathcal{H}_{0}}\alpha_{j}}{1\vee\sum_{j\leq t}\delta_{j}}\approx\operatorname{FDP}(t):=\frac{\sum_{j\leq t,j\in\mathcal{H}_{0}}\mathbb{I}\{p_{j}\leq\alpha_{j}\}}{1\vee\sum_{j\leq t}\delta_{j}},$ (1)
where $\mathcal{H}_{0}$ is the index set of null hypotheses. To ensure that $\operatorname{FDR}(t)\leq\alpha$, it suffices to enforce $\widehat{\operatorname{FDP}}_{\text{LORD}}(t)=\frac{\sum_{j\leq t}\alpha_{j}}{R(t)\vee 1}\leq\alpha.$ The corresponding testing level of LORD++ is
$\alpha_{t}^{\text{LORD++}}=\gamma_{t}s_{0}+(\alpha-s_{0})\gamma_{t-\tau_{1}}\mathbb{I}\{\tau_{1}<t\}+\alpha\sum_{j:\tau_{j}<t,\tau_{j}\neq\tau_{1}}\gamma_{t-\tau_{j}},$
where $\{\gamma_{t}\}_{t=1}^{\infty}$ is a given infinite non-increasing sequence of positive constants that sums to one, $\tau_{j}$ is the time of the $j$-th rejection, and $s_{0}>0$ is the pre-specified initial wealth.
Building on this, some adaptive versions of LORD++ were proposed subsequently, such as SAFFRON [Ramdas et al., 2018] and ADDIS [Tian and Ramdas, 2019]. Specifically, SAFFRON introduces a user-specified parameter $\lambda\in[0,1]$ to estimate the proportion of true nulls in the online testing stream, and obtains a less conservative estimated FDP as
$\widehat{\operatorname{FDP}}_{\text{SAFFRON}}(t)=\frac{\sum_{j\leq t}\alpha_{j}\frac{\mathbb{I}\{p_{j}>\lambda\}}{(1-\lambda)}}{1\vee\sum_{j\leq t}\delta_{j}}.$ (2)
These GAI‐based procedures guarantee online mFDR control under the conditional super‐uniformity assumption for null pp‐values:
@@
where $\mathcal{G}_{t}=\sigma(\delta_{1:t})$ and each threshold $\alpha_{t}=f_{t}(\delta_{1:t-1})$ is measurable with respect to past decisions. Here $\delta_{1:t}:=(\delta_{1},\dots,\delta_{t})$. Furthermore, if all null pp‐values are independent of other pp‐values, these procedures also ensure strict FDR control.
Note that the inequality (i) in (1) is tight if and only if all past hypotheses are null. This conservative step is one of the major sources of the gap between the realized FDR of LORD++ and the target level $\alpha$. When feedback becomes available over time, it allows us to sharpen the FDP estimator and accordingly reduce the resulting slack.